# Sina Rismanchian, Yasaman Razeghi, Sameer Singh, Shayan Doroudi

### v1

@@ 1 Introduction
Large Multimodal Models (LMMs) have the potential to handle tasks that combine visual, linguistic, and reasoning abilities, previously achievable only by humans. Indeed, LMMs such as GPT4-V [52] and Gemini 1.5 flash [46, 17] have been shown to be state of the art models in solving multi-modal tasks such as visual question answering [18, 28], visual mathematical questions [30], chart question answering [33], etc. Despite these successes, there remains the question of how LMMs perform in tasks that intertwine visual reasoning and programming knowledge. That is, given an image of a geometric pattern (and/or a verbal description of the pattern) can LMMs generate code that would be able to procedurally generate that pattern? Indeed, Bubeck et al. [6] showed that the large language model (LLM) with no visual training data was able to create a unicorn in TikZ—the LaTeX-based graphics drawing library. This feat amazed many and provoked many discussions on the intelligence of large language models—but how general is this ability?
@@
In this work, we introduce TurtleBench, a set of manually crafted image/text to code tasks in turtle geometry [37, 1] to evaluate the abilities of these models to combine visual pattern recognition, mathematical reasoning, Python programming, and abstract geometrical reasoning. To ensure the visual inputs and the programming language remain straightforward, TurtleBench harnesses turtle geometry, a concept widely recognized for its effectiveness in introducing programming concepts to children within the K-12 education system. In turtle geometry, a turtle acts as a programmable object that navigates the screen, drawing as it goes and turning at specified angles, to create simple visual patterns. The primary objective within this framework is to generate code capable of producing simple visual inputs. These visual inputs consist of basic geometric shapes, and the programming syntax required is intentionally limited and straightforward. An example of such a task is presented in the left side of Figure 1. As illustrated, the input image is the shape of a simple square and the corresponding code only uses two simple turtle functions (forward and right) along with a simple for loop. This simplicity makes TurtleBench an effective benchmark for evaluating the capabilities of Large Multimodal Models (LMMs). To reflect different real-world use cases of an LMM in the domain of Turtle and also cover the broad range of underlying reasoning abilities, TurtleBench includes 260 tasks with a variety of types and modalities. The different types and input/output modalities are presented in Figure 1. We define two types of tasks scratch and tweak in TurtleBench. Scratch tasks challenge models to generate Python code for a specified shape using the Turtle library based on inputs that could be an image of the shape, a textual description, or both. This subset evaluates model’s proficiency in recognizing patterns within the shape and accurately translating these into executable code. Conversely, tweak tasks are designed to probe deeper into a model’s understanding, examining its ability to comprehend the implications of described modifications to shapes—such as connecting midpoints (example in Figure 1)—and their representation in the image. Here, models are provided with a base shape and are instructed to create the desired alteration in shape. Instructions for these modifications may be provided visually or textually. To simplify, in a subset of the tweak tasks, we adopt a code editing approach, supplying the original shape’s code and directing the model to edit this code to generate the target shape.
@@
We conduct an evaluation of leading LMMs on TurtleBench code generation and code editing tasks, utilizing zero-shot and visual chain-of-thought approach [41] across text-only, image-only, and mixed (text and image) input modalities. Our findings reveal that these models generally perform poorly across all setups and variety of tasks and modalities. Both GPT-4o and Gemini 1.5 Flash struggle with TurtleBench tasks, failing to solve more than 75% of them. Our results show that performance improves when tasks are presented as text rather than images, suggesting that integrating visual and linguistic information, especially for visual pattern recognition, requires further refinement. When tested with a custom library mimicking Python Turtle but using different command names, models showed a significant performance drop, revealing difficulties in generalizing visual reasoning to unfamiliar syntax. Even when allowed to choose their own programming language, models consistently failed to generate correct code, indicating broader challenges in translating visual instructions into functional programming outputs. These findings show that our benchmark poses a significant challenge for LMMs, offering key insights into their limitations. Our evaluation highlights gaps in integrating visual reasoning with programming and raises important questions for future research to address.
@@
In turtle geometry [1], a turtle is a programmable object on the screen that leaves a trace while moving. As illustrated in Figure 1 left side, we see an example of how creating a simple geometric shape, like a square, involves the turtle moving forward and executing turns four times. It is a powerful, intuitive tool that enables novice learners to start learning programming by creating aesthetically beautiful artifacts. Although Turtle programming nowadays is used more as a tool to foster computational thinking, it can be used to teach geometry and mathematical reasoning [31, 9] as it enables learners to explore and learn geometrical relationships between shapes. The intuitive nature and learnability of turtle geometry, along with its ability to generate patterns of diverse complexities, make it a compelling concept upon which to base a benchmark for LMMs. In the following, we describe our benchmark in detail.
@@ 3.1 Models
In the following section, we evaluate TurtleBench using two SOTA LMMs, GPT-4o and Gemini 1.5 Flash and also an open sourced model, namely Llava-1.5-13B [27] employing greedy decoding in our evaluations. We evaluated two other open models, namely Qwen-VL-Max [3] and CogVLM [47] on a subset of tasks in TurtleBench. However, CogVLM and Qwen are not successful in producing a syntactically correct Python Turtle piece of code even for the simplest tasks, therefore we limited our benchmark evaluation to models mentioned above.
@@ 3.2 Prompting
We use two types of prompting in our experiments, 1) basic, where we simply prompt the the model (c.f. Appendix A.2) to do our tasks. , and 2) Chain-of-Thought (CoT) prompting [49], which has shown to be an effective prompting technique in eliciting reasoning in these models. Specifically, we use a more detailed version of CoT prompting that is tailored to LMMs, namely v-CoT, recently proposed by [41]. The v-CoT approach is inspired by m-CoT [54], which shows higher performance compared to it. This prompting has been shown to improve LMMs’ performance on visual tasks that involved reasoning, such as ARC [8]. This prompt, instructs the model to first extract all the relevant information in the image needed for answering the problem and then to reason step by step based on the information extracted. The specific prompt we used in our experiments is in Appendix A.2
@@ 4 Results
Results on the performance of the models are reported in percentage, where in each experiment we evaluate the performance of select models on all instances of the TurtleBench with test@1 method [10] where the model generates only one piece of code for each instance. We assign a binary value to the success/failure of each task. We then run each experiment on a model five different times and we report the average percentage of accumulative success.
@@ 4.1 Models perform poorly on TurtleBench
We initially examine the performance of the GPT-4o, Gemini 1.5 Flash and Llava-1.5-13B models on the comprehensive TurtleBench dataset. The findings, detailed in Table  1, reveal a notably poor performance across the tasks in TurtleBench, with a peak accuracy of 20% achieved by GPT-4o in the code editing tasks, facilitated by Chain of Thought (CoT) prompting. In the scratch tasks, which represent the simplest problem type within the dataset, GPT-4o’s success rate was just 19%, underscoring the substantial challenges and complexities these tasks pose to the current models. A comparison between CoT and basic prompting within Table  1 illustrates that CoT prompting outperforms basic prompting on the same models, aligning with previous work that indicates CoT enhances models’ reasoning abilities [54]. However, despite employing CoT, the task remains far from being solved. Examples of model output in different subsets of the task are provided in Figures 2 and 10.
As previous work suggests that in-context learning by providing examples of the task at hand can significantly improve models’ performance in domain adaptation [5], we evaluated the performance of GPT-4o and Gemini 1.5 Flash on Scratch code generation with 4-shot CoT prompting where we first provided four pairs of shape-code to the model. Yet, we did not find any major improvement in the performance of the models.
@@ 4.3 Assessing Model Proficiency Across Programming Languages
To answer the question, we run an ablation study on GPT-4o as our best-performing model in the main tasks. We allow it to generate code using any library, language, or similar tools it deems appropriate, such as Matplotlib, TikZ, etc., without restricting it to the Python Turtle library. The prompt for this subset of tasks is presented in Appendix A.2.4. We manually evaluate the GPT-4o output for this task. Despite this freedom, we observe no significant improvement in performance. The model chooses Matplotlib for 50% of the tasks and offers pseudocode for 2%, with the remainder reverting to Python Turtle, even though we do not specify Python Turtle in the prompts. Notably, it avoids using TikZ, despite its mention in the prompt and proven capabilities in prior work to produce TikZ code [6, 4]. We further isolate this observation by obligating the model to produce code using Matplotlib (refraining from generating pseudocode and Python Turtle) and in this experiment as well, we do not see a major improvement in the model’s results. This outcome underscores a deeper issue than syntax familiarity: the models’ fundamental challenge is accurately interpreting visual input and applying this understanding to generate corresponding programming code.
@@
To distinctly assess the models’ proficiency in interpreting textual versus visual information, we conducted an evaluation focusing on their ability to reason about the relationship between provided code and corresponding images in Tweak code edit tasks. In this setup, models are given a base shape along with the Python Turtle code that generated it. Subsequently, they are prompted to adjust the given code to modify the shape according to specified instructions. These instructions are delivered in two forms: 1) (I + T) as natural language descriptions, e.g., ’connect the midpoints of each side to the midpoints of its adjacent sides,’ and 2) (I + I) as images explicitly showcasing the desired outcome. Examples of these subsets are in Figure 1. Our comparison of model performance across these two modalities reveals a huge decline in accuracy when instructions were provided visually rather than textually (Table 1). This outcome suggests a disparity in the models’ ability to process visual versus textual instructions, revealing that their reasoning abilities may not align closely with human-like understanding. The assumption that directly viewing the desired outcome simplifies the task contrasts sharply with our findings, highlighting a reliance on textual interpretation for reasoning and a notable limitation in pure visual reasoning capabilities within these models [39].
@@ 4.5 Vision component contributes poorly
One of the questions regarding LMMs’ abilities in visual abstraction and understanding tasks is the extent the incorporation of the visual component has enhanced their abilities in reasoning [34].
In resonance with what Mitchell et al. [34] found, here we also found that the vision component contributes poorly to fostering the models’ visual reasoning abilities, at least in the domain of TurtleBench. Specifically, we annotated 27 (21%) Scratch code generation tasks and provided clear descriptions for each in plain text (Textual descriptions were validated by two human annotators, one in the research team and the other recruited as a volunteer). The remaining shapes were too complex to describe without ambiguity in plain text. Then, we compared the three modes of presenting the task, image only, text only, and the blend of an image and its textual description (I, T, and I + T, respectively in Table 1). Interestingly, for both GPT-4o and Gemini 1.5 Flash, the model performed worse when the task was presented only in the image, compared to the other modes. This phenomenon is counterintuitive as for humans, perceiving the images should be easier than first reading a description, imagining it, and then writing a code for it. Additionally, as presented in Table 1 the blend of image and text only slightly improved GPT-4o’s performance (from 38% to 40%). These two findings show that there is still much room for improvement especially in the visual components of LMMs.
@@ 5.1 Large Multi-modal Models
Recent advancements in foundational multimodal models have marked a significant stride towards developing generalist AI systems capable of understanding and integrating information across different modalities to solve tasks without the need for task-specific fine-tuning. Among these models are closed source models such as Gemini 1.5 Flash [45], GPT-4o [36], and open source models as LLaVA-1.5 [27], Mini-GPT4 [55], InstructBLIP [13] and CogVLM [48]. The versatility and multimodal understanding exhibited by these foundational multimodal models have positioned them as prime candidates for applications such as AI software engineers or programming tutors for children. Our work evaluates the efficacy of these popular models on image/text-to-code tasks, measuring their potential in vision/programming context.
@@ 5.2 Probabilistic Program Induction
Recent work in Bayesian cognitive science has modeled various aspects of cognition and learning as probabilistic program induction [24, 23, 40, 15, 50, 19]. This has involved both modeling human cognition as program induction as well as designing machine learning algorithms that can generate programs for various tasks, including the kind of turtle geometry task we study here. Ellis et al. [15] developed the DreamCoder algorithm which can learn to induce programs by using self-supervision to incrementally build up a library of programs and train a neural network to search to find the best program for a given task. They created a dataset of 160 turtle programming tasks. In contrast to our approach, where we assess the performance of out-of-the-box LMMs, DreamCoder is trained on a training set of images (i.e., half of the dataset). However, it is interesting that the algorithm is trained in an unsupervised fashion; that is, DreamCoder never receives the code used to generate the images and learns that from experience. Wong et al. [50] extended this work by developing an algorithm (LAPS) that can induce programs given both the task and linguistic annotations for the task. They used a dataset of 311 turtle graphics with greater complexity than the original DreamCoder dataset. While their dataset includes linguistic annotations, their dataset does not include tweak tasks like in TurtleBench. Additionally, their tasks often include arbitrary aspects (for example, a gap with unspecified distance between two shapes) that makes evaluation hard; in our tasks, the positional relationships between shapes should be easy to infer exactly and hence we can evaluate models by comparing exactly with ground truth shapes. Moreover, neither of these datasets have been framed as a benchmark for visual program induction and have not been considered for evaluating LMMs. Perhaps the approach closest to our work is by Grand et al. [19], who combined LLMs with a symbolic program induction algorithm and evaluated the performance of their model (LILO) on the turtle geometry task using the aforementioned dataset. Averaged over several runs, the performance of the best versions of these approaches on the turtle geometry task is as follows: 43% for DreamCoder, 82% for LAPS, 49% for LILO, and 32% for a LLM solver. These results seem to suggest that probabilistic programming approaches (such as LAPS) can greatly outperform LMMs on visual programming tasks. We note that the performance of the LLM solver (32%) is comparable to the performance of GPT-4o on our text-only input (37%; see Table 1). Future work could assess the performance of probabilistic program induction methods like LAPS on TurtleBench.

### v2

@@ Abstract
TurtleBench: A Visual Programming Benchmark in Turtle Geometry
Sina Rismanchian, Yasaman Razeghi, Sameer Singh, Shayan Doroudi University of California, Irvine {srismanc,yrazeghi,sameer,doroudis}@uci.edu
@@ 1 Introduction
Large Multimodal Models (LMMs) have the potential to handle tasks that combine visual, linguistic, and reasoning abilities, previously achievable only by humans. Indeed, LMMs such as GPT4-V (Yang et al., 2023) and Gemini 1.5 flash (Team et al., 2023b; Fu et al., 2023) have been shown to be state-of-the-art models in solving multi-modal tasks such as visual question answering (Goyal et al., 2017; Liu et al., 2024), coding for visual multimodal questions (Li et al., 2024a), visual mathematical questions (Lu et al., 2023), chart question answering (Masry et al., 2022), etc. Despite these successes, there remains the question of how LMMs perform in tasks that intertwine visual reasoning and programming knowledge. That is, given an image of a geometric pattern (and/or a verbal description of the pattern) can LMMs generate code that would be able to procedurally generate that pattern? Indeed, Bubeck et al. (2023) showed that the large language model (LLM) with no visual training data was able to create a unicorn in TikZ—the LaTeX-based graphics drawing library. This feat amazed many and provoked many discussions on the intelligence of large language models—but how general is this ability?
@@
In this work, we introduce TurtleBench, a set of manually crafted image/text to code tasks in turtle geometry (Papert, 1972; Abelson and diSessa, 1986) to evaluate the abilities of these models to combine visual pattern recognition, mathematical reasoning, Python programming, and abstract geometrical reasoning. To ensure the visual inputs and the programming language remain straightforward, TurtleBench harnesses turtle geometry, a concept widely recognized for its effectiveness in introducing programming concepts to children within the K-12 education system. In turtle geometry, a turtle acts as a programmable object that navigates the screen, drawing as it goes and turning at specified angles, to create simple visual patterns. The primary objective within this framework is to generate code capable of producing simple visual inputs. These visual inputs consist of basic geometric shapes, and the programming syntax required is intentionally limited and straightforward. An example of such a task is presented in the left side of Figure 1. As illustrated, the input image is the shape of a simple square and the corresponding code only uses two simple turtle functions (forward and right) along with a simple for loop. This simplicity makes TurtleBench an effective benchmark for evaluating the capabilities of Large Multimodal Models (LMMs). To reflect different real-world use cases of an LMM in the domain of Turtle and also cover the broad range of underlying reasoning abilities, TurtleBench includes 260 tasks with a variety of types and modalities. The different types and input/output modalities are presented in Figure 1. We define two types of tasks scratch and tweak in TurtleBench. Scratch tasks challenge models to generate Python code for a specified shape using the Turtle library based on inputs that could be an image of the shape, a textual description, or both. This subset evaluates model’s proficiency in recognizing patterns within the shape and accurately translating these into executable code. Conversely, tweak tasks are designed to probe deeper into a model’s understanding, examining its ability to comprehend the implications of described modifications to shapes—such as connecting midpoints (example in Figure 1)—and their representation in the image. Here, models are provided with a base shape and are instructed to create the desired alteration in shape. Instructions for these modifications may be provided visually or textually. To simplify, in a subset of the tweak tasks, we adopt a code editing approach, supplying the original shape’s code and directing the model to edit this code to generate the target shape.
@@
We conduct an evaluation of leading LMMs on TurtleBench code generation and code editing tasks, utilizing zero-shot and visual chain-of-thought approach (Singh et al., 2023) across text-only, image-only, and mixed (text and image) input modalities. Our findings reveal that these models generally perform poorly across all setups and variety of tasks and modalities. Both GPT-4o and Gemini 1.5 Flash struggle with TurtleBench tasks, failing to solve more than 75% of them. Our results show that performance improves when tasks are presented as text rather than images, suggesting that integrating visual and linguistic information, especially for visual pattern recognition, requires further refinement. When tested with a custom library mimicking Python Turtle but using different command names, models showed a significant performance drop, revealing difficulties in generalizing visual reasoning to unfamiliar syntax. Even when allowed to choose their own programming language, models consistently failed to generate correct code, indicating broader challenges in translating visual instructions into functional programming outputs. These findings show that our benchmark poses a significant challenge for LMMs, offering key insights into their limitations. Our evaluation highlights gaps in integrating visual reasoning with programming and raises important questions for future research to address.
@@
In turtle geometry (Abelson and diSessa, 1986), a turtle is a programmable object on the screen that leaves a trace while moving. As illustrated in Figure 1 left side, we see an example of how creating a simple geometric shape, like a square, involves the turtle moving forward and executing turns four times. It is a powerful, intuitive tool that enables novice learners to start learning programming by creating aesthetically beautiful artifacts. Although Turtle programming nowadays is used more as a tool to foster computational thinking, it can be used to teach geometry and mathematical reasoning Marji (2014); Clements and Sarama (2013) as it enables learners to explore and learn geometrical relationships between shapes. The intuitive nature and learnability of turtle geometry, along with its ability to generate patterns of diverse complexities, make it a compelling concept upon which to base a benchmark for LMMs. In the following, we describe our benchmark in detail.
@@ 3.1 Models
In the following section, we evaluate TurtleBench using two SOTA LMMs, GPT-4o and Gemini 1.5 Flash and also an open sourced model, namely Llava-1.5-13B (Liu et al., 2023) employing greedy decoding in our evaluations. We evaluated two other open models, namely Qwen-VL-Max (Bai et al., 2023) and CogVLM (Wang et al., 2023) on a subset of tasks in TurtleBench. However, CogVLM and Qwen are not successful in producing a syntactically correct Python Turtle piece of code even for the simplest tasks, therefore we limited our benchmark evaluation to models mentioned above.
@@ 3.2 Prompting
We use two types of prompting in our experiments, 1) basic, where we simply prompt the the model (c.f. Appendix A.2) to do our tasks. , and 2) Chain-of-Thought (CoT) prompting (Wei et al., 2022), which has shown to be an effective prompting technique in eliciting reasoning in these models. Specifically, we use a more detailed version of CoT prompting that is tailored to LMMs, namely v-CoT, recently proposed by Singh et al. (2023). The v-CoT approach is inspired by m-CoT (Zhang et al., 2023), which shows higher performance compared to it. This prompting has been shown to improve LMMs’ performance on visual tasks that involved reasoning, such as ARC (Chollet, 2019). This prompt, instructs the model to first extract all the relevant information in the image needed for answering the problem and then to reason step by step based on the information extracted. The specific prompt we used in our experiments is in Appendix A.2
@@ 4 Results
Results on the performance of the models are reported in percentage, where in each experiment we evaluate the performance of select models on all instances of the TurtleBench with test@1 method Cobbe et al. (2021) where the model generates only one piece of code for each instance. We assign a binary value to the success/failure of each task. We then run each experiment on a model five different times and we report the average percentage of accumulative success.
@@ 4.1 Models perform poorly on TurtleBench
We initially examine the performance of the GPT-4o, Gemini 1.5 Flash and Llava-1.5-13B models on the comprehensive TurtleBench dataset. The findings, detailed in Table  1, reveal a notably poor performance across the tasks in TurtleBench, with a peak accuracy of 20% achieved by GPT-4o in the code editing tasks, facilitated by Chain of Thought (CoT) prompting. In the scratch tasks, which represent the simplest problem type within the dataset, GPT-4o’s success rate was just 19%, underscoring the substantial challenges and complexities these tasks pose to the current models. A comparison between CoT and basic prompting within Table  1 illustrates that CoT prompting outperforms basic prompting on the same models, aligning with previous work that indicates CoT enhances models’ reasoning abilities (Zhang et al., 2023). However, despite employing CoT, the task remains far from being solved. Examples of model output in different subsets of the task are provided in Figures 2 and 10.
As previous work suggests that in-context learning by providing examples of the task at hand can significantly improve models’ performance in domain adaptation Brown (2020), we evaluated the performance of GPT-4o and Gemini 1.5 Flash on Scratch code generation with 4-shot CoT prompting where we first provided four pairs of shape-code to the model. Yet, we did not find any major improvement in the performance of the models.
@@ 4.3 Assessing Model Proficiency Across Programming Languages
To answer the question, we run an ablation study on GPT-4o as our best-performing model in the main tasks. We allow it to generate code using any library, language, or similar tools it deems appropriate, such as Matplotlib, TikZ, etc., without restricting it to the Python Turtle library. The prompt for this subset of tasks is presented in Appendix A.2.4. We manually evaluate the GPT-4o output for this task. Despite this freedom, we observe no significant improvement in performance. The model chooses Matplotlib for 50% of the tasks and offers pseudocode for 2%, with the remainder reverting to Python Turtle, even though we do not specify Python Turtle in the prompts. Notably, it avoids using TikZ, despite its mention in the prompt and proven capabilities in prior work to produce TikZ code (Bubeck et al., 2023; Belouadi et al., 2023). We further isolate this observation by obligating the model to produce code using Matplotlib (refraining from generating pseudocode and Python Turtle) and in this experiment as well, we do not see a major improvement in the model’s results. This outcome underscores a deeper issue than syntax familiarity: the models’ fundamental challenge is accurately interpreting visual input and applying this understanding to generate corresponding programming code.
@@
To distinctly assess the models’ proficiency in interpreting textual versus visual information, we conducted an evaluation focusing on their ability to reason about the relationship between provided code and corresponding images in Tweak code edit tasks. In this setup, models are given a base shape along with the Python Turtle code that generated it. Subsequently, they are prompted to adjust the given code to modify the shape according to specified instructions. These instructions are delivered in two forms: 1) (I + T) as natural language descriptions, e.g., ’connect the midpoints of each side to the midpoints of its adjacent sides,’ and 2) (I + I) as images explicitly showcasing the desired outcome. Examples of these subsets are in Figure 1. Our comparison of model performance across these two modalities reveals a huge decline in accuracy when instructions were provided visually rather than textually (Table 1). This outcome suggests a disparity in the models’ ability to process visual versus textual instructions, revealing that their reasoning abilities may not align closely with human-like understanding. The assumption that directly viewing the desired outcome simplifies the task contrasts sharply with our findings, highlighting a reliance on textual interpretation for reasoning and a notable limitation in pure visual reasoning capabilities within these models Roberts and Roberts (2024).
@@ 4.5 Vision component contributes poorly
One of the questions regarding LMMs’ abilities in visual abstraction and understanding tasks is the extent the incorporation of the visual component has enhanced their abilities in reasoning (Mitchell et al., 2023).
In resonance with what Mitchell et al. (2023) found, here we also found that the vision component contributes poorly to fostering the models’ visual reasoning abilities, at least in the domain of TurtleBench. Specifically, we annotated 27 (21%) Scratch code generation tasks and provided clear descriptions for each in plain text (Textual descriptions were validated by two human annotators, one in the research team and the other recruited as a volunteer). The remaining shapes were too complex to describe without ambiguity in plain text. Then, we compared the three modes of presenting the task, image only, text only, and the blend of an image and its textual description (I, T, and I + T, respectively in Table 1). Interestingly, for both GPT-4o and Gemini 1.5 Flash, the model performed worse when the task was presented only in the image, compared to the other modes. This phenomenon is counterintuitive as for humans, perceiving the images should be easier than first reading a description, imagining it, and then writing a code for it. Additionally, as presented in Table 1 the blend of image and text only slightly improved GPT-4o’s performance (from 38% to 40%). These two findings show that there is still much room for improvement especially in the visual components of LMMs.
@@ 5.1 Large Multi-modal Models
Recent advancements in foundational multimodal models have marked a significant stride towards developing generalist AI systems capable of understanding and integrating information across different modalities to solve tasks without the need for task-specific fine-tuning. Among these models are closed source models such as Gemini 1.5 Flash (Team et al., 2023a), GPT-4o (OpenAI et al., 2024), and open source models as LLaVA-1.5 (Liu et al., 2023), Mini-GPT4 (Zhu et al., 2023), InstructBLIP (Dai et al., 2023) and CogVLM (Wang et al., 2024). The versatility and multimodal understanding exhibited by these foundational multimodal models have positioned them as prime candidates for applications such as AI software engineers or programming tutors for children. Our work evaluates the efficacy of these popular models on image/text-to-code tasks, measuring their potential in vision/programming context.
@@ 5.2 Probabilistic Program Induction
Recent work in Bayesian cognitive science has modeled various aspects of cognition and learning as probabilistic program induction (Lake et al., 2015; Lake and Piantadosi, 2020; Rule et al., 2020; Ellis et al., 2023; Wong et al., 2021; Grand et al., 2023). This has involved both modeling human cognition as program induction as well as designing machine learning algorithms that can generate programs for various tasks, including the kind of turtle geometry task we study here. Ellis et al. (2023) developed the DreamCoder algorithm which can learn to induce programs by using self-supervision to incrementally build up a library of programs and train a neural network to search to find the best program for a given task. They created a dataset of 160 turtle programming tasks. In contrast to our approach, where we assess the performance of out-of-the-box LMMs, DreamCoder is trained on a training set of images (i.e., half of the dataset). However, it is interesting that the algorithm is trained in an unsupervised fashion; that is, DreamCoder never receives the code used to generate the images and learns that from experience. Wong et al. (2021) extended this work by developing an algorithm (LAPS) that can induce programs given both the task and linguistic annotations for the task. They used a dataset of 311 turtle graphics with greater complexity than the original DreamCoder dataset. While their dataset includes linguistic annotations, their dataset does not include tweak tasks like in TurtleBench. Additionally, their tasks often include arbitrary aspects (for example, a gap with unspecified distance between two shapes) that makes evaluation hard; in our tasks, the positional relationships between shapes should be easy to infer exactly and hence we can evaluate models by comparing exactly with ground truth shapes. Moreover, neither of these datasets have been framed as a benchmark for visual program induction and have not been considered for evaluating LMMs. Perhaps the approach closest to our work is by Grand et al. (2023), who combined LLMs with a symbolic program induction algorithm and evaluated the performance of their model (LILO) on the turtle geometry task using the aforementioned dataset. Averaged over several runs, the performance of the best versions of these approaches on the turtle geometry task is as follows: 43% for DreamCoder, 82% for LAPS, 49% for LILO, and 32% for a LLM solver. These results seem to suggest that probabilistic programming approaches (such as LAPS) can greatly outperform LMMs on visual programming tasks. We note that the performance of the LLM solver (32%) is comparable to the performance of GPT-4o on our text-only input (37%; see Table 1). Future work could assess the performance of probabilistic program induction methods like LAPS on TurtleBench.